Joe spent $15.75 on 21 pieces of fruit to make a salad. Bananas cost $.50 each, oranges cost $.075 each, and papayas cost $1.25 each. Joe used twice as many oranges as bananas. How many of each fruit did he buy?


1
Expert's answer
2021-02-25T01:31:14-0500

"x - bananas\\
y - oranges\\
z - papayas\\"

"\begin{cases}
 x + y + z = 21\\
 0.5x + 0.75y + 1.25z = 15.75\\
 y = 2x
\end{cases}" "\begin{cases}
 x + 2x + z = 21\\
 0.5x + 0.75\cdot2x + 1.25z = 15.75\\
 y = 2x
\end{cases}"

"\begin{cases}
 3x + z = 21\\
 2x + 1.25z = 15.75\\
 y = 2x
\end{cases}\\" "\begin{cases}
 z = 21 - 3x\\
 2x + 1.25z = 15.75\\
 y = 2x
\end{cases}" "\begin{cases}
 z = 21 - 3x\\
 2x + 1.25\cdot(21 - 3x) = 15.75\\
 y = 2x
\end{cases}
\begin{cases}
 z = 21 - 3x\\
 2x + 26.25 - 3.75x = 15.75\\
 y = 2x
\end{cases}
\begin{cases}
 z = 21 - 3x\\
 -1.75x = 15.75 - 26.25\\
 y = 2x
\end{cases}\\
\begin{cases}
 z = 21 - 3x\\
 1.75x = 10.5\\
 y = 2x
\end{cases}
\begin{cases}
 z = 21 - 3x\\
 x = 6\\
 y = 2x
\end{cases}
\begin{cases}
 z = 21 - 18\\
 x = 6\\
 y = 12
\end{cases}
\begin{cases}
 z = 3\\
 x = 6\\
 y = 12
\end{cases}\\\\"

Answer: 6 bananas, 12 oranges, 3 papayas